# Doubly Robust Alignment for Large Language Models

### v1

@@ 4 Double Robust Preference Evaluation and Optimization
$\widehat{p}_{\textrm{DR}}(\pi)=\mathbb{E}_{(X,Y^{(1)},Y^{(2)},Z)\sim\mathcal{D}}\psi(X,Y^{(1)},Y^{(2)},Z;\pi,\widehat{\pi}_{\textrm{ref}},\widehat{g}).$ (9)
Similar to the DR estimator in the bandit setting [Dudík et al., , 2014], (9) is reduced to the DM estimator when setting $\widehat{g}$ to zero, and the IS estimator when setting the IS ratio $\pi/\widehat{\pi}_{\textrm{ref}}$ to zero. However, as shown in (8), a key different from those bandit estimators is that in our pairwise comparison setting, each data tuple is used twice – as $(X,Y^{(1)},Y^{(2)},Z)$ and $(X,Y^{(2)},Y^{(1)},1-Z)$ – in constructing the estimating function. This effectively reduces the variance of the resulting estimator. As a result, we will formally show in Section 5 that our DR estimator is semi-parametrically efficient. Additionally, we will establish the consistency of (9) when either $\widehat{g}$ or $\widehat{\pi}_{\textrm{ref}}$ is correctly specified.
Preference optimization.
For preference optimization, our goal is to identify the optimal policy that maximizes the average total preference $p^{*}(\pi)$. Under the BT model assumption, it is immediate to see that the argmax is equivalent to $\pi^{*}$ defined in (2). Given the proposed DR estimator, we estimate the optimal policy by solving
$\widehat{\pi}=\arg\max_{\pi\in\Pi}\Big{\{}\widehat{p}_{\textrm{DR}}(\pi)-\beta\mathbb{E}_{X\sim\mathcal{D}}{D}_{\mathrm{KL}}[\pi(\cdot\mid X)\,\|\,\widehat{\pi}_{\mathrm{ref}}(\cdot\mid X)]\Big{\}}.$ (10)
We refer to (10) as DRPO, short for doubly robust preference optimization. Theoretically, we will show in Section 5 that our estimated policy $\widehat{\pi}$ achieves a smaller regret bound than PPO- and DPO-type algorithms when the BT assumption holds. Practically, we implement three refinements to stabilize the training: (i) clipping the IS ratio to avoid extremely large IS ratio; (ii) designing a pseudo objective function to enable Monte Carlo sampling from the target policy during optimization; (iii) adopting the KL divergence measure from the group relative policy optimization [Shao et al., , 2024] for variance reduction. Details are relegated to Appendix B to save space.
@@ 5 Theoretical Analysis
We begin with a summary of our theories; Figure 2 outlines the roadmap. Our theories are concerned with (i) the MSE of our preference evaluation estimator $\widehat{p}_{\textrm{DR}}(\pi)$ (see (9)), and (ii) the regret bounds of $\widehat{\pi}$ (see (10)) computed by the proposed preference optimization algorithm. Specifically, Theorem 5.5 provides a finite sample upper bound for the MSE of $\widehat{p}_{\textrm{DR}}(\pi)$, which in turn yields its double robustness (Corollary 5.6) and semi-parametric efficiency (Corollary 5.7). Meanwhile, Theorem 5.8 upper bounds the difference in total preference between the optimal in-class policy and $\widehat{\pi}$, without assuming the BT model holds. It reveals the double robustness property of our preference optimization algorithm (Corollary 5.9). When the BT model holds, Theorem 5.10 further upper bounds the suboptimal gap of $\widehat{\pi}$, demonstrating that it general achieves smaller gaps than PPO- and DPO-based algorithms.
@@
Assumption 5.1 (Coverage).
Both $\pi/\pi_{\textrm{ref}}$ and $\pi/\widehat{\pi}_{\textrm{ref}}$ are upper bounded by $\epsilon^{-1}$ for some constant $\epsilon>0$.
Assumption 5.2 (Boundedness).
@@
Assumption 5.4 (Model complexity).
$\Pi$ belongs to the Vapnik–Chervonenkis (VC) type class [Chernozhukov et al., , 2014, Definition 2.1] with a finite VC index $v>0$.
We remark that similar coverage, boundedness and realizability assumptions are commonly imposed in the OPE and RL literature [see e.g., Chen and Jiang, , 2019, Fan et al., , 2020, Uehara et al., , 2022]. The VC-class condition is also frequently assumed in statistics and machine learning [see e.g., Van Der Vaart et al., , 1996, Shalev-Shwartz and Ben-David, , 2014].
MSE of $\widehat{p}_{\textrm{DR}}(\pi)$.
We next study the statistical properties of the proposed preference estimator $\widehat{p}_{\textrm{DR}}(\pi)$. Without loss of generality, we also assume both $\widehat{\pi}_{\textrm{ref}}$ and $\widehat{g}$ (or $\widehat{r}$, in the case where the BT model holds) are obtained from external models independent of $\mathcal{D}$. This condition is mild. Even when such external models are not available and $\widehat{\pi}_{\textrm{ref}}$ and $\widehat{g}$ are learned internally from $\mathcal{D}$, independence can be preserved using sample-splitting and cross-fitting [Chernozhukov et al., , 2018].
Theorem 5.5 (MSE).
Under Assumption 5.1, with $n$ data tuples, the semi-parametric efficiency bound (SEB) for estimating $p^{*}(\pi)$ is given by $n^{-1}\textrm{Var}(\psi(X,Y^{(1)},Y^{(2)},Z;\pi,\pi_{\textrm{ref}},g^{*}))$. Additionally, the MSE of our $\widehat{p}_{\textrm{DR}}(\pi)$ equals
@@
The first part of Theorem 5.5 establishes the SEB – the smallest-possible MSE that one can hope for estimating $p^{*}(\pi)$. The second part upper bounds the excess MSE of our estimator over SEB. Specifically, this excess MSE consists of three parts: the first two are excess variance terms arising from estimation errors in the reference policy and the preference model, while the third is a bias term introduced by these estimation errors. Notably, both variance terms and SEB scale as $O(n^{-1})$. Consequently, when either $\widehat{\pi}_{\textup{ref}}$ or $\widehat{g}$ is correctly specified, the MSE of $\widehat{p}_{\textrm{DR}}(\pi)$ coverages to zero as the sample $n$ approaches to infinity. This establishes the double robustness property of our estimator, which we state below.
Corollary 5.6 (Doubly robust evaluation).
@@
Under Assumption 5.1, when either $\widehat{\pi}_{\textup{ref}}$ or $\widehat{g}$ is correctly specified, the MSE of $\widehat{p}_{\textrm{DR}}(\pi)$ decays to zero as $n$ approaches to infinity.
@@
We next consider the case where both $\widehat{\pi}_{\textup{ref}}$ and $\widehat{g}$ are “approximately” correct in that both root MSEs $\|\widehat{\pi}_{\textup{ref}}/\pi_{\textup{ref}}-1\|$ and $\|\widehat{g}-g^{*}\|$ decay to zero as $n\to\infty$. Since SEB is of the order $O(n^{-1})$, the first two variance terms in (11) decay to zero at a much faster rate than SEB. Meanwhile, when the product $\|\widehat{\pi}_{\textup{ref}}/\pi_{\textup{ref}}-1\|\|\widehat{g}-g^{*}\|=o(n^{-1/2})$, the last bias term in (11) becomes negligible compared to SEB as well. Together, these conditions imply that the MSE of $\widehat{p}_{\textrm{DR}}(\pi)$ asymptotically matches the SEB, which establishes the semi-parametric efficiency of our estimator. We also remark that conditions similar to $\|\widehat{\pi}_{\textup{ref}}/\pi_{\textup{ref}}-1\|\|\widehat{g}-g^{*}\|=o(n^{-1/2})$ are widely assumed in the literature [see e.g., Chernozhukov et al., , 2018, Farrell et al., , 2021, Kallus and Uehara, , 2022].
Corollary 5.7 (Semi-parametric efficiency).
Under Assumption 5.1, when both $\|\frac{\widehat{\pi}_{\textup{ref}}}{\pi_{\textup{ref}}}-1\|$ and $\|\widehat{g}-g^{*}\|$ decay to zero as $n\to\infty$, and their product is $o(n^{-1/2})$, then MSE($\widehat{p}_{\textrm{DR}}(\pi)$)/SEB $\to 1$ as $n\to\infty$.
Regret of $\widehat{\pi}$.
@@
Next, we derive the statistical properties of the proposed policy $\widehat{\pi}$. When the BT model assumption is violated, we measure the regret of a given policy $\pi$ using the gap between the total preference of the best in-class policy and that of $\pi$, i.e., $\textrm{Reg}(\pi)=\sup_{\pi^{\prime}\in\Pi}p^{*}(\pi^{\prime})-p^{*}(\pi)$. By definition, a smaller regret indicates a better policy.
Theorem 5.8 (Regret).
Under Assumptions 5.1 (assuming it holds for any $\pi\in\Pi$) and 5.4, then
$\textrm{Reg}(\widehat{\pi})=O\Big{(}\beta+\sqrt{\frac{v}{n}}+\frac{v}{n}+\|\widehat{\pi}_{\textup{ref}}/\pi_{\textup{ref}}-1\|\|\widehat{g}-g^{*}\|\Big{)}.$ (12)
@@
It can be seen from (12) that the regret bound depends on several factors: (i) it decays with the sample size $n$; (ii) it increases with the regularization parameter $\beta$ in the KL divergence penalty; (iii) it increases with $v$, which measures the complexity of the policy class; (iv) it decreases with the estimating error of the reference policy and the preference model. Crucially, the last dependence appears as the product $\|\widehat{\pi}_{\textup{ref}}/\pi_{\textup{ref}}-1\|\|\widehat{g}-g^{*}\|$, which enables us to establish the double robustness property in the context of preference optimization.
Corollary 5.9 (Doubly robust optimization).
@@
Suppose $\beta\to 0$ as $n\to 0$. Under the conditions in Theorem 5.8, when either $\widehat{\pi}_{\textup{ref}}$ or $\widehat{g}$ is correctly specified, the regret of $\widehat{\pi}$ decays to zero as $n\to\infty$.
Finally, we restrict our attention to the ideal setting where the BT model holds and upper bound the suboptimality gap, defined as the difference in the expected reward between the optimal policy $\pi^{*}$ and our $\widehat{\pi}$, i.e., $J(\pi^{*})-J(\widehat{\pi})$.
Theorem 5.10 (Suboptimality gap).
@@
Suppose the BT model assumption in (1) holds. Under Assumptions 5.2, 5.3, and the conditions in Theorem 5.8, the suboptimality gap of $\widehat{\pi}$ is upper bounded by
@@
$O\Big{(}\exp(-\bar{c}\beta^{-1})+\frac{1}{\beta}\sqrt{\frac{v}{n}}+\|\widehat{\pi}_{\textup{ref}}/\pi_{\textup{ref}}-1\|\Big{)},$ (15)
@@
According to (12) and (14) that, by using a sufficiently small $\beta$, the suboptimality gaps of PPO-based and our algorithms are of the order $O(n^{-1/2}+\|\widehat{r}-r^{*}\|)$ and $O(n^{-1/2}+\|\widehat{\pi}_{\textup{ref}}/\pi_{\textup{ref}}-1\|\|\widehat{r}-r^{*}\|)$, respectively. As for DPO-based algorithms, For DPO-based algorithms, setting $\beta=\bar{c}^{-1}C\log n$ for some constant $C>0$ makes the first term in (15) of order $O(n^{-C})$, which can be made arbitrarily small with a sufficiently large $C$. The second term remains of order $O(n^{-1/2})$ up to a logarithmic factor, yielding an overall suboptimality gap of $O(n^{-1/2}\log n+\|\widehat{\pi}_{\textup{ref}}/\pi_{\textup{ref}}-1\|)$. Consequently, our algorithm’s suboptimality gap is more robust to estimation errors in the reference policy and preference model, as these errors influence our bound only through their product. To the contrary, for PPO- and DPO-based algorithms, these errors affect their suboptimality bounds in the first order. In particular, when these errors converge to zero at a rate of $O(n^{-c})$ for some $20<c<1/$ 2, our algorithm achieves strictly smaller suboptimality bounds than both DPO- and PPO-based algorithms.
@@ 6 Experiments
In this section, we first use the IMDb dataset [Maas et al., , 2011] to empirically validate the double robustness property of our preference estimator $\widehat{p}_{\textrm{DR}}$ (Equation 9) established in Corollary 5.6. We next compare the proposed preference optimization algorithm (Equation 10) against baseline approaches on the Too Long; Didn’t Read [TL;DR, Völske et al., , 2017] and Anthropic Helpful and Harmless [HH, Bai et al., , 2022] datasets.
6.1 Preference Evaluation
@@
Task and objective.
@@
We consider the controlled sentiment generation task which aims to produce positive movie reviews using the IMDb dataset. We first apply SFT to the EleutherAI/gpt-neo-125m base model [Black et al., , 2021], which serves as the reference policy for response generation. The generated responses are then annotated using a pre-trained sentiment classifier to produce preference labels. Using these synthetic data, we train an optimal policy via DPO. Our objective in this section is to evaluate the total preference of this DPO-trained policy over the SFT-based reference policy. Its oracle value, computed via Monte Carlo, is 0.681. Additional details on data generation and model training are provided in Appendix C.1.
Evaluation and result.
@@
To empirically assess the double robustness property, we evaluate four variants of our preference estimator, each with either the preference model and/or the reference policy correctly specified or misspecified. To misspecify the preference model, we set $\widehat{g}$ to a uniformly random value in $[0,1]$. To misspecify the reference policy, we use the unfine-tuned EleutherAI/gpt-neo-125m base model. Figure 3 displays the MSEs (solid lines on left panel) and their associated 95% confidence intervals (shaded areas) of the four estimators across different sample sizes, averaged over 500 simulations. It can be seen that the estimator with both models misspecified (red line) exhibits a significantly larger MSE than the other three and shows minimal improvement beyond 800 samples. To the contrary, when either the preference model or the reference policy is correctly specified (yellow and green lines), the MSE is substantially reduced with a moderately large sample size. This aligns with the double robustness property. Meanwhile, the estimator with both correctly specified models (blue line) achieves the lowest MSE (being very close to zero with 1500 data tuples), supporting its semiparametric efficiency.
6.2 Preference Optimization
Tasks and objective.
This section considers two tasks: summarization and human dialogue. First, for summarization, we use the TL;DR dataset with preference annotations from [Stiennon et al., , 2020] (lengthy Reddit posts as prompts, behavioral SFT models’ summaries as responses) to fine-tune LLMs for concise, informative summaries. The SFT and reward models for this task are from cleanrl [Huang et al., , 2024], with the SFT model trained on a similar but different dataset. Second, for human dialogue, the HH dataset (human queries as prompts) is used to align LLMs for helpful responses. Its SFT and reward models are trained using TRL [von Werra et al., , 2020] framework due to the unavailability of pre-existing ones. For each task, a reward-based BT preference model (using the same reward model for PPO training) and a general preference model [Zhang et al., 2024d, ] are adopted to serve as $\widehat{g}$ (donating as DRPO-BT and DRPO-GPM). Refer to more details of the implementation and baseline model training in Appendix C.2.
@@ 6.2 Preference Optimization
Figure 4: Pairwise win rate matrices between different methods across two datasets. Left: TL;DR dataset. Right: HH dataset. Each entry indicates how often the row method outperforms the column method; higher values denote better performance.
Evaluation and result.
Given the absence of ground-truth preference or reward models, we follow prior work [Rafailov et al., , 2023, Wu et al., 2024a, , Ye et al., , 2025] and use GPT-4o-mini to assess the quality of generated responses (details in Appendix C.2). Win rates at default temperature 1.0 are presented in Figure 4, with results at other temperatures presented in Appendix D. In summarization, both DRPO-BT and DRPO-GPM largely outperform DPO and PPO. For this task, we use an external reference policy model for $\widehat{\pi}_{\mathrm{ref}}$, which is likely misspecified due to being trained on a different dataset. This likely contributes to DPO’s poor performance. However, despite relying on the same reference policy, the superior performance of our algorithms highlights their robustness to reference policy misspecification. In human dialogue, DRPO-GPM demonstrates the best performance, whereas DRPO-BT outperforms PPO and achieves comparable performance to DPO. The relatively weaker performance of PPO suggests potential misspecification in the reward model for this task. Despite using the same reward model, DRPO-BT effectively mitigates this challenge and achieves a win rate of 57% against PPO, demonstrating its robustness. Finally, it is worth highlighting that the hyperparameters for our method are not extensively tuned, unlike the baselines – particularly PPO, for which the best-performing version is reported – further demonstrating the robustness of our proposal.

### v2

@@ 4 Double Robust Preference Evaluation and Optimization
[…]
Similar to the DR estimator in the bandit setting [112], (9) is reduced to the IS estimator when setting $\widehat{g}$ to zero, and the DM estimator when setting the IS ratio $\pi/\widehat{\pi}_{\textrm{ref}}$ to zero. However, as shown in (8), a key different from those bandit estimators is that in our pairwise comparison setting, each data tuple is used twice – as $(X,Y^{(1)},Y^{(2)},Z)$ and $(X,Y^{(2)},Y^{(1)},1-Z)$ – in constructing the estimating function. This effectively reduces the variance of the resulting estimator. As a result, we will formally show in Section 5 that our DR estimator is semi-parametrically efficient. Additionally, we will establish the consistency of (9) when either $\widehat{g}$ or $\widehat{\pi}_{\textrm{ref}}$ is correctly specified.
Preference optimization. For preference optimization, our goal is to identify the optimal policy that maximizes the average total preference $p^{*}(\pi)$. Under the BT model assumption, it is immediate to see that the argmax is equivalent to $\pi^{*}$ defined in (2). Given the proposed DR estimator, we estimate the optimal policy by solving
$\widehat{\pi}=\arg\max_{\pi\in\Pi}\Big\{\widehat{p}_{\textrm{DR}}(\pi)-\beta\mathbb{E}_{X\sim\mathcal{D}}{D}_{\mathrm{KL}}[\pi(\bullet\mid X)\,\|\,\widehat{\pi}_{\mathrm{ref}}(\bullet\mid X)]\Big\}.$ (10)
We refer to (10) as DRPO, short for doubly robust preference optimization. Theoretically, we will show in Section 5 that our estimated policy $\widehat{\pi}$ achieves a smaller suboptimality gap bound than PPO- and DPO-based algorithms when the BT assumption holds. Practically, we implement three refinements to stabilize the training: (i) clipping the IS ratio to avoid extremely large IS ratio; (ii) designing a pseudo objective function to enable Monte Carlo sampling from the target policy during optimization; (iii) adopting the KL divergence measure from the group relative policy optimization [43] for variance reduction. Details are relegated to Appendix B to save space.
@@ 5 Theoretical Analysis
We begin with a summary of our theories; Figure 2 outlines the roadmap. Our theories are concerned with (i) the MSE of our preference evaluation estimator $\widehat{p}_{\textrm{DR}}(\pi)$ (see (9)), and (ii) the performance gap bounds of $\widehat{\pi}$ (see (10)) computed by the proposed preference optimization algorithm. Specifically, Theorem 2 provides a finite sample upper bound for the MSE of $\widehat{p}_{\textrm{DR}}(\pi)$, which in turn yields its double robustness (Corollary 3) and semi-parametric efficiency (Corollary 4). Meanwhile, Theorem 5 upper bounds the difference in total preference between the optimal in-class policy and $\widehat{\pi}$, without assuming the BT model holds. It reveals the double robustness property of our preference optimization algorithm (Corollary 6). When the BT model holds, Theorem 7 further upper bounds the suboptimal gap of $\widehat{\pi}$, demonstrating that it general achieves smaller gaps than PPO- and DPO-based algorithms.
@@
Assumption 1 (Coverage).
$\pi/\pi_{\textrm{ref}}$ and $\pi/\widehat{\pi}_{\textrm{ref}}$ are upper bounded by $\epsilon^{-1}$ for some constant $\epsilon>0$.
Assumption 2 (Boundedness).
@@
Assumption 4 (Model complexity).
$\Pi$ belongs to the Vapnik–Chervonenkis (VC) type class [144, Definition 2.1] with a finite VC index $v>0$.
We remark that similar coverage, boundedness and realizability assumptions are commonly imposed in the OPE and RL literature [see e.g., 145, 146, 143]. The VC-class condition is also frequently assumed in statistics and machine learning [see e.g., 147, 148].
MSE of $\widehat{p}_{\textrm{DR}}(\pi)$. We next study the statistical properties of the proposed preference estimator $\widehat{p}_{\textrm{DR}}(\pi)$. Without loss of generality, we also assume both $\widehat{\pi}_{\textrm{ref}}$ and $\widehat{g}$ (or $\widehat{r}$, in the case where the BT model holds) are obtained from external models independent of $\mathcal{D}$. This condition is mild. Even when such external models are not available and $\widehat{\pi}_{\textrm{ref}}$ and $\widehat{g}$ are learned internally from $\mathcal{D}$, independence can be preserved using sample-splitting and cross-fitting [93].
Theorem 2 (MSE).
Under Assumption 1, with $n$ data tuples, the semi-parametric efficiency bound (SEB) for estimating $p^{*}(\pi)$ is given by $n^{-1}\textrm{Var}(\psi(X,Y^{(1)},Y^{(2)},Z;\pi,\pi_{\textrm{ref}},g^{*}))$. Additionally, the MSE of our $\widehat{p}_{\textrm{DR}}(\pi)$ equals
@@
The first part of Theorem 2 establishes the SEB – the smallest-possible MSE that one can hope for estimating $p^{*}(\pi)$. The second part upper bounds the excess MSE of our estimator over SEB. Specifically, this excess MSE consists of three parts: the first two are excess variance terms arising from estimation errors in the reference policy and the preference model, while the third is a bias term introduced by these estimation errors. Notably, (i) it can be shown that SEB scales as $O(n^{-1})$; (ii) the two variance terms decrease to zero as the sample size $n$ approaches infinity; (iii) the bias term is a product of the MSEs of $\widehat{\pi}_{\textup{ref}}$ and $\widehat{g}$. Consequently, when either $\widehat{\pi}_{\textup{ref}}$ or $\widehat{g}$ is correctly specified, the MSE of $\widehat{p}_{\textrm{DR}}(\pi)$ converges to zero as $n$ approaches to infinity. This establishes the double robustness property of our estimator, which we state below.
Corollary 3 (Doubly robust evaluation).
@@
Under Assumption 1, when either $\widehat{\pi}_{\textup{ref}}$ or $\widehat{g}$ is correctly specified, the MSE of $\widehat{p}_{\textrm{DR}}(\pi)$ decays to zero as $n$ approaches to infinity.
@@
We next consider the case where both $\widehat{\pi}_{\textup{ref}}$ and $\widehat{g}$ are “approximately” correct in that both root MSEs $\|\widehat{\pi}_{\textup{ref}}/\pi_{\textup{ref}}-1\|$ and $\|\widehat{g}-g^{*}\|$ decay to zero as $n\to\infty$. Since SEB is of the order $O(n^{-1})$, the first two variance terms in (11) decay to zero at a even faster rate than SEB. Meanwhile, when the product $\|\widehat{\pi}_{\textup{ref}}/\pi_{\textup{ref}}-1\|\|\widehat{g}-g^{*}\|=o(n^{-1/2})$, the last bias term in (11) becomes negligible compared to SEB as well. Together, these conditions imply that the MSE of $\widehat{p}_{\textrm{DR}}(\pi)$ asymptotically matches the SEB, which establishes the semi-parametric efficiency of our estimator. We also remark that conditions similar to $\|\widehat{\pi}_{\textup{ref}}/\pi_{\textup{ref}}-1\|\|\widehat{g}-g^{*}\|=o(n^{-1/2})$ are widely assumed in the literature [see e.g., 93, 103, 82, 149].
Corollary 4 (Semi-parametric efficiency).
[…]
@@
Regret of $\widehat{\pi}$. Next, we derive the statistical properties of the proposed policy $\widehat{\pi}$. When the BT model assumption is violated, we measure the performance gap of a given policy $\pi$ using the gap between the total preference of the best in-class policy and that of $\pi$, i.e., $\textrm{Gap}(\pi)=\sup_{\pi^{\prime}\in\Pi}p^{*}(\pi^{\prime})-p^{*}(\pi)$. By definition, a smaller performance gap indicates a better policy.
Theorem 5 (Performance gap).
Under Assumptions 1 (assuming it holds for any $\pi\in\Pi$) and 4, then
$\textrm{Gap}(\widehat{\pi})=O\Big(\beta+\sqrt{\frac{v}{n}}+\frac{v}{n}+\|\widehat{\pi}_{\textup{ref}}/\pi_{\textup{ref}}-1\|\|\widehat{g}-g^{*}\|\Big).$ (12)
@@
It can be seen from (12) that the performance gap depends on several factors: (i) it decays with the sample size $n$; (ii) it increases with the regularization parameter $\beta$ in the KL divergence penalty; (iii) it increases with $v$, which measures the complexity of the policy class; (iv) it decreases with the estimating error of the reference policy and the preference model. Crucially, the last dependence appears as the product $\|\widehat{\pi}_{\textup{ref}}/\pi_{\textup{ref}}-1\|\|\widehat{g}-g^{*}\|$, which enables us to establish the double robustness property in the context of preference optimization.
Corollary 6 (Doubly robust optimization).
@@
Suppose $\beta\to 0$ as $n\to 0$. Under the conditions in Theorem 5, when either $\widehat{\pi}_{\textup{ref}}$ or $\widehat{g}$ is correctly specified, Gap($\widehat{\pi}$) decays to zero as $n\to\infty$.
[…]
Theorem 7 (Suboptimality gap).
@@
Suppose the BT model assumption in (1) holds. Under Assumptions 2, 3, and the conditions in Theorem 5, the suboptimality gap of $\widehat{\pi}$ is upper bounded by
@@
$O\Big(\exp(-\bar{c}\beta^{-1})+\beta^{-1}\sqrt{\frac{v}{n}}+\|\widehat{\pi}_{\textup{ref}}/\pi_{\textup{ref}}-1\|\Big),$ (15)
@@
According to (13) and (14) that, by using a sufficiently small $\beta$, the suboptimality gaps of PPO-based and our algorithms are of the order $O(n^{-1/2}+\|\widehat{r}-r^{*}\|)$ and $O(n^{-1/2}+\|\widehat{\pi}_{\textup{ref}}/\pi_{\textup{ref}}-1\|\|\widehat{r}-r^{*}\|)$, respectively. As for DPO-based algorithms, setting $\beta=\bar{c}^{-1}C\log n$ for some constant $C>0$ makes the first term in (15) of order $O(n^{-C})$, which can be made arbitrarily small with a sufficiently large $C$. The second term remains of order $O(n^{-1/2})$ up to a logarithmic factor, yielding an overall suboptimality gap of $O(n^{-1/2}\log n+\|\widehat{\pi}_{\textup{ref}}/\pi_{\textup{ref}}-1\|)$. Consequently, our algorithm’s suboptimality gap is more robust to estimation errors in the reference policy and preference model, as these errors influence our bound only through their product. To the contrary, for PPO- and DPO-based algorithms, these errors affect their suboptimality bounds in the first order. In particular, when these errors converge to zero at a rate of $O(n^{-c})$ for some $0<c<1/2$, our algorithm achieves strictly smaller suboptimality bounds than both DPO- and PPO-based algorithms.
To conclude this section, we make two remarks. First, a key novelty of our analysis lies in the derivation of DPO’s sub-optimality bounds without relying on linearity assumptions. While there is extensive literature on DPO-based algorithms, their sub-optimality gaps are relatively underexplored. Some recent works derive such bounds under strong linear assumptions, which simplify the analysis by allowing the sub-optimality gap to be expressed directly in terms of parameter estimation error [150]. In contrast, our analysis proceeds without such linear assumptions, which makes the derivation much more challenging. Second, Theorem 7 establishes upper bounds on the sub-optimality gaps, and we discuss the tightness of these bounds in Appendix A.7.
@@ 6 Experiments
In this section, we first use the IMDb dataset [151] to empirically validate the double robustness property of our preference estimator $\widehat{p}_{\textrm{DR}}$ (Equation 9) established in Corollary 3. We next compare the proposed preference optimization algorithm (Equation 10) against baseline approaches on the Too Long; Didn’t Read [TL;DR, 152] and Anthropic Helpful and Harmless [HH, 8] datasets.
@@
These datasets are particularly suited for studying preference and/or reference model misspecification. Specifically: (i) TL;DR illustrates reference model misspecification – we use the SFT model trained by CleanRL [153], which was learned on a filtered subset of the data, leading to a misspecified reference policy; (ii) HH illustrates preference model misspecification, as prior works suggest this dataset contains unmodeled pairwise noise beyond BT [53, 54]; (iii) IMDb illustrates both types of misspecification, since it is synthetic dataset where we have access to the ground-truth preference and reference models.
@@
Preference Evaluation. We consider the controlled sentiment generation task which aims to produce positive movie reviews using the IMDb dataset. We first apply SFT to the EleutherAI base model [154], which serves as the reference policy for response generation. The generated responses are then annotated using a pre-trained sentiment classifier to produce preference labels. Using these synthetic data, we train an optimal policy via DPO. Our objective in this section is to evaluate the total preference of this DPO-trained policy over the SFT-based reference policy. Its oracle value, computed via Monte Carlo, is 0.681. Additional details on data generation and model training are provided in Appendix C.1.
@@
To empirically assess the double robustness property, we evaluate four variants of our preference estimator, each with either the preference model and/or the reference policy correctly specified or misspecified. To misspecify the preference model, we set $\widehat{g}$ to a uniformly random value in $[0,1]$. To misspecify the reference policy, we use the unfine-tuned EleutherAI base model. Figure 3 displays the MSEs (solid lines on left panel) and their associated 95% confidence intervals (shaded areas) of the four estimators across different sample sizes, averaged over 500 simulations. It can be seen that the estimator with both models misspecified (red line) exhibits a significantly larger MSE than the other three and shows minimal improvement beyond 800 samples. To the contrary, when either the preference model or the reference policy is correctly specified (yellow and green lines), the MSE is substantially reduced with a moderately large sample size. This aligns with the double robustness property. Meanwhile, the estimator with both correctly specified models (blue line) achieves the lowest MSE (being very close to zero with 1500 data tuples), supporting its semiparametric efficiency.
@@
Figure 4: Pairwise win rate matrices between different methods across two datasets. Left: TL;DR. Right: $H$. Each entry indicates how often the row method outperforms the column method.
Preference Optimization. This section considers two tasks: summarization and human dialogue. First, for summarization, we use the TL;DR dataset, where lengthy Reddit posts serve as prompts and preference annotations are from Stiennon et al. [140], to fine-tune LLMs for concise, informative summaries. Both SFT and reward models for this task are obtained from CleanRL. Second, for human dialogue, the HH dataset (human queries as prompts) is used to align LLMs for helpful responses. Since the original SFT and reward models are unavailable, we train them ourselves using the TRL framework [155]. For each task, a reward-based BT preference model (using the same reward model for PPO training) and a general preference model [19] are adopted to serve as $\widehat{g}$ (donating as DRPO-BT and DRPO-GPM). Refer to more details of the implementation and baseline model training in Appendix C.2.
We compare our DRPO against nine baseline fine-tuning algorithms, including the standard PPO and DPO, and seven variants of DPO: (i) Dr. DPO [54]; (ii) rDPO [53]; (iii) cDPO [156]; (iv) CPO [157]; (v) ORPO [158]; (vi) IPO [56]; (vii) RSO [48]. Given the absence of ground-truth preference and reward models, we adopt two evaluation strategies. The first strategy uses in-distribution data. Specifically, for both TL;DR and HH, one portion of the dataset is used to fine-tune the LLMs, while the remaining portion is used to generate responses for evaluation. Following prior works [26, 41, 40], we employ GPT-4o-based annotator to compare the quality of responses produced by two LLMs (details in Appendix C.2). Win rates – the percentage of cases in which one LLM’s response is preferred over another – are reported at the default temperature of 1.0 in Figure 4 and Table 3, with results at other temperatures provided in Appendix D. The second strategy uses the out-of-distribution data provided via the AlpacaEval 2.0 benchmark [159], which covers a broad collection of human-written instructions designed for general-purpose tasks. Pairwise comparisons are conducted using a GPT-4-Turbo-based annotator. Since summarization is a domain-specific task, we apply the out-of-distribution evaluation only to human dialogue (Table 3).
In summarization, both DRPO-BT and DRPO-GPM substantially outperform PPO, DPO (see the left panel of Figure 4), and DPO’s variants (Table 3). As mentioned earlier, the reference policy in this dataset is misspecified, likely contributing to the weaker performance of DPO and its variants. Despite trained on the same misspecified reference policy, the superior performance of DRPO highlights its robustness to such misspecification. In human dialogue, DRPO-GPM demonstrates the best in-distribution performance, whereas DRPO-BT outperforms PPO and achieves comparable performance to DPO (see the right panel of Figure 4). The poor performance of PPO partly supports the potential misspecification of BT in this task. Despite using the same preference model, DRPO-BT achieves a win rate of 57% against PPO, demonstrating its robustness. As for out-of-distribution evaluation, DRPO performs comparably to robust DPO variants (cDPO, rDPO and Dr. DPO) while attaining higher win rates than other variants (Table 3). As discussed earlier, the HH dataset likely contains pairwise noise, which these robust variants are explicitly designed to handle, whereas DRPO employs a preference model that does not account for such noise. If DRPO were to adopt the same noise-aware preference model used in these methods, its performance would likely improve further.
Table 2: Win rates of DRPO (using BT as the preference model) compared to various baseline algorithms on TL;DR. Higher win rates indicate better performance of DRPO over the baseline algorithm.
Baseline Model Win Rate (%)
DRPO vs Dr. DPO 72.5
DRPO vs rDPO 65.0
DRPO vs cDPO 63.5
DRPO vs CPO 90.0
DRPO vs ORPO 57.5
DRPO vs IPO 98.5
DRPO vs RSO 69.5
Table 3: Win rates of different algorithms compared to SFT on $H$. “LC Win Rate” denotes the length-controlled win rate. DRPO uses GPM as preference model.
Model LC Win Rate (%) Win Rate (%)
Dr. DPO 92.16 90.93
rDPO 86.89 85.71
cDPO 85.05 84.28
CPO 73.59 71.28
ORPO 75.92 53.91
IPO 78.29 78.88
RSO 80.62 79.50
DRPO 86.38 84.84